# revtex4-2Repair the float

### v1

@@ 2 Simplest Example
Compare [21, 263-265]. Consider $V$ as the span of $\text{i}/2\sigma_{x},\text{i}/2\sigma_{z}$, the controls parametrized as $c\left(t\right)=\text{i}/2\left(c_{x}\left(t\right)\sigma_{x}+c_{z}\left(t\right)\sigma_{z}\right)$. The sub-Riemannian geodesics have time-dependent coefficients $c\left(t\right)$ of the form
@@
$\begin{split}&U\stackrel{{\scriptstyle!}}{{=}}C\left(t\right)\\ &=\exp\left[\frac{\text{i}}{2}t\left(\cos\phi_{0}\sigma_{x}+\sin\phi_{0}\sigma_{z}+\beta\sigma_{y}\right)\right]\exp\left[-\frac{\text{i}}{2}t\beta\sigma_{y}\right].\end{split}$ (10)
Given the rotation to implement, the parameters of the geodesic can be found. For most cases, this involves solving a nonlinear equation for $\beta$. In (Fig 2) there are examples for a geodesic with a unitary $\text{e}^{\text{i}\alpha\sigma_{y}}$.
@@
Arclength for Rotation in $y$ DirectionEstimates of the sub-Riemannian $\text{d}_{CC}$ Distance
$d_{CC}$$C_{1,2}\cdot d_{\infty}$$L,d_{R}$
Figure 4: We compare the different Bounds for a $Y$-rotation in the $\text{SU}(2)$ with $X$ and $Z$ as horizontal directions. Riemannian distance or arclength and scaling for the Euler angles (Blue). The Sub-Riemannian distance that is important to our error bound (Black). Bounds with $d_{\infty}$ given by the box-ball theorem (Red).
@@ V Conclusion and Outlook
Several questions remain open. First, can an analogous error minimization framework be developed for the correlated regime of MBQC, where operations are not independent and inter-site correlations play a critical role?—Both the prsesent work and the work [20] that explains how to operate in the correlated regime, are based off of [12]. The latter provides a first scaling relation such as Eq. (3), but under the limitations of applicability (a) to only the uncorrelated regime, and (b) to only single generators. Ref. [20] lifts restriction (a), and the present work lifts restriction (b). It is desirable to have a scaling result that lifts both restrictions simultaneously.
@@ Acknowledgements.
LH acknowledges helpful discussions with Tom Tischel, Paul Herringer, Wolfram Bauer, Hendrik P. Nautrup and Vir B. Bulchandani. This work was supported by the Alexander von Humboldt foundation.

### v2

@@ 2 Simplest Example
Compare [3, 263-265]. Consider $V$ as the span of $\mathrm{i}/2\sigma_{x},\mathrm{i}/2\sigma_{z}$, the control parametrized as $c\left(t\right)=\mathrm{i}/2\left(c_{x}\left(t\right)\sigma_{x}+c_{z}\left(t\right)\sigma_{z}\right)$. The sub-Riemannian geodesics have time-dependent coefficients $c\left(t\right)$ of the form
@@
$\begin{split}&U\stackrel{{\scriptstyle!}}{{=}}C\left(T\right)\\ &=\exp\left[\frac{\mathrm{i}}{2}T\left(\cos\phi_{0}\sigma_{x}+\sin\phi_{0}\sigma_{z}+\beta\sigma_{y}\right)\right]\exp\left[-\frac{\mathrm{i}}{2}T\beta\sigma_{y}\right].\end{split}$ (11)
Given the rotation to implement, the parameters of the geodesic can be found. For most cases, this involves solving a nonlinear equation for $\beta$. In (Fig 2) there are examples for a geodesic with a unitary $\mathrm{e}^{\mathrm{i}\alpha/2\sigma_{y}}$.
@@
Arclength for Rotation in $y$ DirectionBounds of the sub-Riemannian $\text{d}_{CC}$ Distance
$d_{\mathrm{CC}}$$C_{1,2}\cdot d_{\infty}$$L,d_{R}$
Figure 4: We compare the different Bounds for a $Y$-rotation in the $\mathrm{SU}(2)$ with $X$ and $Z$ as horizontal directions. Riemannian distance or arclength and scaling for the Euler angles (Blue). The Sub-Riemannian distance that is important to our error bound (Black). Bounds with $d_{\infty}$ given by the box-ball theorem (Red).
@@ V Conclusion and Outlook
Several questions remain open. First, can an analogous error minimization framework be developed for the correlated regime of MBQC  [1, 2], where operations are not independent and inter-site correlations play a critical role?— Ref.  [31] provided the first scaling relation of the type of Eq. (3), but under the limitations of applicability (a) to only the uncorrelated regime, and (b) to only single generators. Ref. [1] lifts restriction (a), and the present work lifts restriction (b). It is desirable to have a scaling result that lifts both restrictions simultaneously.
@@ Acknowledgements.
LH acknowledges helpful discussions with Tom Tischel, Paul Herringer, Wolfram Bauer, Hendrik P. Nautrup and Vir B. Bulchandani. This work was supported by the Alexander von Humboldt Stiftung.
Data Availability Statement
The depicted data as well as the source code are publicly available on GitHub: [17].